# Binary-Report Peer Prediction for Real-Valued Signal Spaces

### v1

@@ 5.2 Gaussian Model
We also derive the exact form of the best response $\hat{\tau}=c(\rho)\tau$ in § D. Now, note that for any $\tau$, $\hat{\tau}=c(\rho)\tau$ for $c(\rho)\in(0,1)$. Under our dynamics, $\dot{\tau}=\hat{\tau}-\tau.$ If $\tau>0$, then $\hat{\tau}<\tau$, while if $\tau<0$, $\hat{\tau}>\tau$. It follows that $\tau=0$ is stable, while by topological necessity the uninformative equilibria are unstable.
While the equilibria and dynamics behave the same in RBTS as in DG for the Gaussian model, we can ask if there are differences in the convergence rate to stable point $\tau=0$. In fact, we can observe that the best response in both settings is a linear function of $\tau$ with coefficients $m_{\textrm{DG}}$, $m_{\textrm{RBTS}}$ respectively (see § D for a formal derivation). It turns out that in any setting of parameters $a$ and $b$, $m_{\textrm{RBTS}}>m_{\textrm{DG}}$, so that if both mechanisms begin with the same starting threshold $\tau(0)$, the stable equilibrium at 00 is reached faster in DG than RBTS. In a peer grading setting, then, graders will more quickly converge to a half-good half-bad quality consensus than RBTS graders, so that the designer may have more time to react and reset expectations in the latter case.
6 Experiments
We now aim to numerically explore existence and dynamics of equilibria across mechanisms in settings other than the noisy Gaussian model. We consider two natural departures from the “nice” properties of Gaussians, skewness and multimodality.
6.1 Skewness
Skewed models of information capture settings where there is some asymmetry in the tasks. To study such settings, we adapt the same noisy signal model as our running Gaussian example, but with a base distribution which is skewed. Formally, each agent $i$ receives a signal $X_{i}$ such that $X_{i}=Z+Z_{i}$, where $Z$ is the skewed Normal distribution with mean 0 and variance 1, and $Z_{i}\sim N(0,1)$. $Z$ has another parameter, $\alpha$, which controls the skewness of the distribution. The magnitude of $\alpha$ increases the magnitude of skewness; moreover, $\alpha<0$ leads to a left-skewed distribution, and $\alpha>0$ a right-skewed distribution. We study how threshold equilibria move as a function of the parameter $\alpha$.
As depicted in Figure 4, we find a unique equilibrium under both OA and DG which shifts toward the longer tail, i.e. the thresholds move from negative to positive as the skewness moves from left to right (and we recover the $\tau=0$ equilibrium in the original Gaussian model when $\alpha=0$). Thus we expect when there is some asymmetry in the distribution of signals, agents in DG will settle at a threshold that balances out the asymmetry in signal mass. While we observe the same high-level dynamics for OA, these equilibria are unstable and so we still expect agents will move toward an uninformative consensus. For DMI, meanwhile, we note that the same equilibria will occur as DG by our theoretical results. Moreover, given the behavior observed under the Gaussian, we expect the stable equilibria in RBTS to share the same pattern as DG in tracking the skewness of the distribution, but at a rate in between OA and DG.
6.2 Multimodality
We now explore the effects that a distribution with multiple peaks has on the equilibria that occur under our signal model. Consider a classic mixed Gaussian setting where each agent’s signal $X_{i}$ is received i.i.d. from one of $K$ Gaussian components, each with mean $\mu_{k}$ and variance $\sigma_{k}^{2}$. There is an underlying parameter $z\sim\text{Categorical}(\pi)$ for $\pi\in\Delta^{K}$ that determines which Gaussian component all the signals come from. For the purposes of our experiments, we fix $\pi$ to be the uniform distribution over the $K$ components. We numerically calculate the functions $F$, $G$, and $Q$, and check for the sufficient and necessary conditions of our theorems, to find equilibria and their stability in this setting. We set the precision of each component to be the same, and then vary this precision parameter to generate bifurcation graphs and study how these equilibria evolve for $K=2$ and $3$ (see § E for $K=4$).
Figure 4: $F$, $G$, and $Q$ over signals $x$ in the two-component Gaussian mixture setting with means $-2,2$ and precision 2.
Figure 5: Bifurcation diagrams for two- and three-component Gaussian mixtures. (Top) Two-component case: OA (left) exhibits three equilibria at high precision, while DG (right) maintains a single stable equilibrium at $\tau=0$. (Bottom) Three-component case: OA (left) has five equilibria, with two stable at $\tau=\pm 1$; DG (right) has three equilibria, with stability reversed from the two-component OA case.
We first observe that based on our experiments, the number of equilibria increases by two each time we add a Gaussian component (Figure 5). But OA begins with three equilibria for large enough precision under the two-component model, while DG begins with one. Since a unique equilibrium in DG is stable, this is a testament to DG being robust to uninformative equilibria. However, for a high enough separation between the two modes, stable equilibria emerge (at $\tau=0$ in the three component case) in OA, and the area of starting conditions that lead to this equilibrium increases. One can view this phenomenon as a reflection of the original binary signal model: with enough separation between the two modes of the richer signal space, we essentially return to a setting where Output Agreement is useful because the “signal” is effectively which mode has been chosen.
We can view equilibria in the snapshot of a specific precision value in Figure 4. Here, the intersection of $G$ with 1/2 corresponds to equilibria in OA, with $F$ corresponds to equilibria in DG (and DMI), and with $Q$ corresponds to equilibria in RBTS. The function $Q$ closely tracks $G$ for high enough precision values. We therefore find that the same stability and number of equilibria occur in RBTS as in DG. However, we note that bifurcation (the increase in equilibria) occurs at lower precision values for RBTS (see § E for visualization). From a design perspective, a lower bifurcation point in RBTS is nice because there is a wider set of signal distributions allowing for some choice over stable thresholds. Meanwhile, our theoretical results show that equilibria restricted to the threshold space in DMI correspond to those in DG, so that we would expect the trends we observe over precision to transfer directly from DG to DMI.
In general, then, our simulations show that multiple threshold equilibria occur across mechanisms with numbers increasing as modes separate, but similar patterns of stability emerge. Moreover, it is even more clear how different behavior emerges under a richer signal model. One could imagine two different real-valued distributions, the first unimodal as in a peer grading setting and the second bimodal as in a data labeling setting, with the same collapsed binary signal model. While previous results on the collapsed space concludes the same behavior would emerge in both settings, our results show that the number and stability of equilibria will be significantly different.
7 Discussion
We study behavior in several peer prediction mechanisms when signals are real-valued and reports are binary. While all these mechanisms have various truthfulness guarantees in the binary signal setting, in our model we find that the notion of truthfulness breaks. Specifically, when agents play according to thresholds which partition the signal space into “high” or “low” regions, under dynamics we find equilibria rarely stay at the initial threshold a mechanism designer sets. Instead, for several mechanisms (Dasgupta–Ghosh, Determinant Mutual Information, and the Robust Bayesian Truth Serum) agents will often stabilize at a different threshold according to the specific payment scheme. Moreover, in Output Agreement we find that uninformative equilibria are often stable, so that in many information structure settings agents will always shift toward an uninformative consensus. We show how such equilibria play out in a natural, noisy Gaussian signal setting and extend our analysis to experiments with multimodal and skewed distributions.
Our results imply that the standard modeling assumptions in peer prediction literature miss important nuances, and these holes can propagate to real changes in how we expect agents to behave in practice. In particular, models of peer prediction have often taken as given an arbitrary joint distribution $P$ over pairs of signals drawn from a finite set. This suggests that we should be able to choose an arbitrary meaning for each signal and then peer prediction should “just work” for the resulting distribution. In contrast, our results show that peer prediction is substantially more inflexible in practice. Whenever the meaning of signals has endogeneity to it there will typically be only a small number of plausible equilibria in this richer space.
Adding flexibility.
We saw in the Gaussian case that a 50/50 report distribution was a natural outcome, which might be unsuitable for some applications. For example in peer grading or norm enforcement [Alechina et al., 2017], one might want the final decision for most tasks to be “accept”. So what should a mechanism designer do? Since peer prediction naturally relies on multiple reports about each task, one option is to use a decision rule, which combines multiple reports on a task in a way that is aware of this limitation. In peer grading an assignment might be satisfactory if any peer grader reports $H$. Doing so would allow more flexibility on the final decisions, e.g. increasing the fraction of satisfactory assignments, even under fixed semantics of $H$ and $L$.
An alternative is to seek to design new families of peer prediction mechanisms that allow control over where the equilibrium thresholds occur. The fact that our characterizations yielded different conditions among our four mechanisms suggests at least some scope for this, although the the fact that in the Gaussiance case these conditions coincided also points to potential limitations. Nevertheless, this seems an interesting direction for future work.
Beyond binary reports.
[…]
Effort.
Studying effort would be natural in our model. For example, one could model effort in the Gaussian model as giving the agent $aZ+bZ_{i}$ where $b$ is decreasing in their exerted effort; in the limit, they obtain the “ground truth” and need only contend with the noisy reports of others. One interesting interplay with our dynamics is that low-effort agents seem most likely to move the threshold, as high-effort agents will have more reason to respect the current one; e.g. when all agents are high-effort, they all coordinate perfectly on the correct report.
@@ 7 Discussion
We have peeled back the peer prediction abstraction by one layer, from binary signals to real-valued ones. This already substantially changed the behavior of mechanisms. Yet even this is quite a stylized abstraction of the information available. Another missing piece is the fact that in many settings some deterministic context, like the content of the essay being graded, is shared by all looking at the same task. From this “shared signal” comes the well-known problem of spurious correlations where in principle participants may choose strategies that depend on inessential details of the context, like the first word of the essay, to correlate their reports without correlating with the ground truth, all while getting optimal payoffs in multi-task mechansisms. While there is work seeking to deal with heterogeneity among agent behaviors [Agarwal et al., 2020], such behavior seems impossible to fully rule out. But even if such pathological equilibra seem unlikely in systems with many well-meaning participants, there is often still ample metadata about tasks such as categories or difficulties that naturally influence participants’ process of translating their actual experience into a “signal.” Designing mechanisms that are robust in the face of such richness is a further challenge.
@@ Acknowledgments
Appendix A Output Agreement
@@ A.2 Dynamics Generalization
[…]
@@
[…]
@@
Appendix B Dasgupta-Ghosh
B.1 Omitted Proofs
Proof of Proposition 3.
Let $\tau^{*}=\infty$. Then for all signals $x$ such that $x\leq\tau^{*}$, $P(\tau^{*};x)=1\geq\Pr[X^{\prime}\leq\infty]$ (since any $X^{\prime}\in\mathbb{R}$ satisfies $X^{\prime}<\tau^{*}$ with probability one.) Meanwhile, Statement 12 is vacuous since no signal $x\in\mathbb{R}$ satisfies $x>\infty$. A similar argument follows for $\tau^{*}=-\infty$: for all signals $x>\tau^{*}$, $P(\tau^{*};x)=0\leq\Pr[X^{\prime}\leq-\infty]$, while Statement 11 is vacuous since no signal $x\in\mathbb{R}$ satisfies $x<-\infty$. ∎
Proof of Theorem 3.
@@ B.1 Omitted Proofs
Proof of Proposition 4.
@@
Proof of Theorem 4.
@@
B.2 Equilibrium Characterization Generalization
@@
Take the function $H(x)=G(x)-F(x)$. Then Condition 2 implies $H(a)=G(a)-F(a)>0$ and $H(b)=G(b)-F(b)<0$. Since $G$ and $F$ are continuous, $H$ is continuous; thus we can apply the Intermediate Value Theorem to conclude there exists a point $\tau^{*}\in[a,b]$ such that $H(\tau^{*})=0$. Then $\tau^{*}$ is an equilibrium by Theorem 3. ∎
@@
First, by Theorem 8, we know at least one equilibrium $\tau^{*}$ exists in the interval $I$. Now assume that $\Pr[X^{\prime}\leq m\mid m]>1/2$. In other words, since $F(m)=1/2$, $G(m)>F(m)$. It follows under Condition 2 that either $m<a$ or $m\in I$. If $m<a$, then by Theorem 8 there exists an equilibrium $\tau^{*}$ in the interval $I$; $\tau^{*}\geq a>m$, so the result follows. Otherwise, consider the case where $m\in I$. Then define $H(x)=G(x)-F(x)$. $H(m)>0$, while under Condition 2, $H(b)<0$. By the Intermediate Value Theorem, since $H$ is continuous, there exists a point $\tau^{*}\in[m,b]$ such that $H(\tau^{*})=0$. By our characterization in Theorem 3, $\tau^{*}$ is an equilibrium.
@@
B.3 Dynamics Generalization
Under reasonable behavior in the tails, we expect existence of a stable nontrivial equilibria, while the uninformative equilibria remain unstable.
@@ B.3 Dynamics Generalization
[…]
@@
We know already from Theorem 8 that at least one equilibrium exists in the interval $I$. Again, we let $H(x)=G(x)-F(x)$. Note first that since $H(a)>0$ and $H(b)<0$, by continuity and differentiability $H$ must cross 0 at some point $\tau_{0}$ with $H^{\prime}(\tau_{0})<0$. Assume not: then $H(x)\geq 0$ for all $x$, which contradicts the fact that $H(x)<0$ for $x\geq b$. It follows by Theorem 4 that $\tau_{0}$ is a stable equilibrium.
@@
    Pick the equilibrium $\tau_{2}$ which is closest to $b$ (this could be the same as the $\tau_{1}$ we chose in the previous step). By Condition 2 and since $H(x)$ is continuous, $H(x)<0$ for $x>\tau_{2}$. It follows taht $H^{\prime}(\tau_{2})\leq 0$. Then $\tau_{2}$ is stable, at least on the right (if $H^{\prime}(\tau_{2})=0$, then $\tau_{2}$ is stable on the right and unstable on the left).
[…]
@@
Appendix C Omitted Proofs for DMI
As discussed in Section 4, the behavior of DMI under joint-task strategies, which allow agents to map all 4 signals to all 4 reports simultaneously, differs between the traditional binary signal model and our real-valued signal model. In particular, Theorem 9 shows that truthfulness is a Bayes Nash among all joint-task strategies for the binary signal model, but below we show that this property fails for the real-valued signal model.
To get some intuition for these results, first consider the binary signal model and an agent who receives signals $H,L,L,L$. From the form of the mechanism, if they report truthfully, their score will be zero deterministically, since $M_{34}$ will have rank 1. Thus, they might consider deviating, say flipping the last $L$ to $H$. It turns out doing so keeps their expected score at zero, since the other agent is equally likely to align or misalign with their reports. The consistency of the other agent’s strategies is crucial for this last fact; given any slight difference in their strategy for tasks 3 and 4, the first agent would deviate from truthfulness.
Thus, given real-valued signals $(10,-10,-10,-1)$, the optimal report is $(H,L,L,H)$, which is not “truthful” for the prescribed threshold $\tau=0$. Hence DMI is no longer truthful in this joint-task sense, even at the equilibrium threshold $\tau=0$.
C.1 Proof of Theorem 9
First, we prove that in the binary signal model, truthfulness is an equilibrium among all joint-task strotegies $\sigma:\{H,L\}^{T}\to\Delta(\{H,L\}^{T})$. Truthfulness remains a best response even when the other agent deviates from truthful, but crucially, still plays a consistent strategy, which here means strategies that (a) map each signal to a distribution over reports, and (b) use the same such map for every task.
Definition 3.
We show a stronger property: truthfulness maximizes the expected payoff of $M_{\textrm{DMI}}$ when all other agents play consistent strategies. As a corollary, truthfulness is a Bayes Nash equilibrium even over the larger space of strategies that allow one to choose all $T$ reports simultaneously after looking at all $T$ signals.

### v2

@@ 6 Discussion
Another missing piece in our model is the fact that in many settings some deterministic context, like the content of the essay being graded, is shared by all looking at the same task. In principle, participants may choose strategies that depend on these inessential details, like the first word of the essay, to correlate their reports, all while getting optimal payoffs in multi-task mechanisms. While there is work dealing with behavior heterogeneity [Agarwal et al., 2020], such behavior seems impossible to fully rule out. Moreover, even if participants are well-meaning, there is often still ample metadata about tasks such as categories or difficulties that naturally influence participants’ process of translating their actual experience into a “signal.”
Studying effort would also be natural in our model. For example, one could model effort in the Gaussian model as giving the agent $aZ+bZ_{i}$ where $b$ is decreasing in their exerted effort. One interesting interplay with our dynamics is that low-effort agents seem most likely to move the threshold, as high-effort agents will have more reason to respect the current one; e.g. when all agents are high-effort, they all coordinate perfectly on the correct report.
CA with other score matrices.
While the tridiagonal matrix $\Delta_{T}$ is a natural score matrix for Correlated Agreement when signals follow the symmetric Gaussian model, using other score matrices could be useful in settings of multimodality or skewness (e.g. rewarding larger clusters of report bins when the distribution is skewed toward them). Future work could therefore include studying how to ideally set the score matrix when the principal has some knowledge about the underlying information structure, as well as better understanding the relationship between behavior under a fixed score matrix and the true correlation structure of signals.
@@ Acknowledgments
Appendix A Omitted Results for Output Agreement
@@ A.2 Dynamics generalization
Let the agent signal structure satisfy Condition 1 for interval $I$, and such that $G$ is continuous and differentiable. Then there exists an equilibrium $\tau^{*}\in I$ which is unstable, while the equilibria $\pm\infty$ are stable.
@@
We know by Condition 1 that $G(x)\neq 1/2$ for $x\notin[a,b]$, so that no equilibria occur outside the interval $I$ other than $\pm\infty$. Since both $\tau_{1}$ and $\tau_{2}$ (or just $\tau_{1}$, if they are the same equilibrium) are unstable (at least to the left for $\tau_{1}$, and right for $\tau_{2}$), it is topologically necessary that the uninformative equilibria $\pm\infty$ are stable. ∎
@@
Appendix B Omitted Results for Dasgupta-Ghosh
B.1 Equilibrium characterization
We begin by formally characterizing equilibria for DG, as in OA. By Equation 23, the condition for a symmetric Bayes-Nash equilibrium can be written as
For a threshold equilibrium, Condition 12 is equivalent to:
$\forall x\leq\tau,P(\tau;x)\geq\pi_{L}$ (13)
$\forall x>\tau,P(\tau;x)\leq\pi_{L}.$ (14)
Now we have the tools to analyze threshold equilibria in DG. We first observe that uninformative strategies $\tau^{*}=\pm\infty$ remain equilibria in the DG mechanism.
Proposition 5.
$\tau^{*}=\pm\infty$ are both always threshold equilibria under DG.
Proof.
Let $\tau^{*}=\infty$. Then for all signals $x$ such that $x\leq\tau^{*}$, $P(\tau^{*};x)=1\geq\Pr[X^{\prime}\leq\infty]$ (since any $X^{\prime}\in\mathbb{R}$ satisfies $X^{\prime}<\tau^{*}$ with probability one.) Meanwhile, Statement 14 is vacuous since no signal $x\in\mathbb{R}$ satisfies $x>\infty$. A similar argument follows for $\tau^{*}=$-$\infty$: for all signals $x>\tau^{*}$, $P(\tau^{*};x)=0\leq\Pr[X^{\prime}\leq-\infty]$, while Statement 13 is vacuous since no signal $x\in\mathbb{R}$ satisfies $x<$-$\infty$. ∎
Theorem 4.
Let finite threshold $\tau$ be given and $P(\tau;x)$ be continuous in $x$. If $\tau$ is a threshold equilibrium under the DG mechanism then $G(\tau)=F(\tau)$. Conversely, if $G(\tau)=F(\tau)$ and either (a) $P(\tau;x)$ is monotone decreasing in x or (b) $P(\tau;x)-\pi_{L}$ has a single crossing of 0 from positive to negative then $\tau$ is a threshold equilibrium under the DG mechanism.
Proof.
@@ B.1 Equilibrium characterization
B.2 Dynamics
We consider stability of equilibria for DG under the dynamics described by Equation (6). Note that as in OA, when everyone else is playing according to $\tau$, there exists a unique best response which is also a threshold strategy.
Proposition 6.
Proof.
@@ B.2 Dynamics
Now, as in OA, we can characterize the stability of equilibria.
Theorem 5.
Proof.
@@
B.3 Gaussian model
Corollary 3.
In the Gaussian model under DG, we have three equilibria at $\tau=0$ and $\pm\infty$.
Proof.
Uniqueness of $\tau=0$ as a finite threshold equilibrium follows. ∎
Now, consider the stability of these equilibria.
Corollary 4.
In the Gaussian model under DG, the threshold equilibrium $\tau=0$ is stable, while the uninformative equilibria $\pm\infty$ are unstable.
Proof.
As depicted in Figure 2, G(x)F(x)G(x)-F(x) is strictly decreasing at 0. Formally, $G^{\prime}(x)-F^{\prime}(x)=c_{G}\phi(c_{G}x)-c_{F}\phi(c_{F}x)$, so $G^{\prime}(0)-F^{\prime}(0)=(c_{G}-c_{F})\phi(0)<0$ (since $c_{G}<c_{F}$ by (15)). Thus, by Theorem 5, the equilibrium at 0 is stable. Since $\tau=0$ is stable, the equilibria $\tau=\pm\infty$ are unstable by topological necessity. ∎
B.4 Equilibrium characterization generalization
@@
Take the function H(x)=G(x)F(x)H(x)=G(x)-F(x). Then Condition 2 implies H(a)=G(a)F(a)>0H(a)=G(a)-F(a)>0 and H(b)=G(b)F(b)<0H(b)=G(b)-F(b)<0. Since $G$ and $F$ are continuous, $H$ is continuous; thus we can apply the Intermediate Value Theorem to conclude there exists a point $\tau^{*}\in[a,b]$ such that $H(\tau^{*})=0$. Then $\tau^{*}$ is an equilibrium by Theorem 4. ∎
@@
First, by Theorem 6, we know at least one equilibrium $\tau^{*}$ exists in the interval $I$. Now assume that $\Pr[X^{\prime}\leq m\mid m]>1/2$. In other words, since $F(m)=1/2$, $G(m)>F(m)$. It follows under Condition 2 that either $m<a$ or $m\in I$. If $m<a$, then by Theorem 6 there exists an equilibrium $\tau^{*}$ in the interval $I$; $\tau^{*}\geq a>m$, so the result follows. Otherwise, consider the case where $m\in I$. Then define H(x)=G(x)F(x)H(x)=G(x)-F(x). $H(m)>0$, while under Condition 2, $H(b)<0$. By the Intermediate Value Theorem, since $H$ is continuous, there exists a point $\tau^{*}\in[m,b]$ such that $H(\tau^{*})=0$. By our characterization in Theorem 4, $\tau^{*}$ is an equilibrium.
@@
B.5 Dynamics generalization
Under reasonable behavior in the tails, we expect existence of a stable, nontrivial equilibrium, while the uninformative equilibria remain unstable.
@@ B.5 Dynamics generalization
Let the agent signal structure satisfy Condition 2 for interval $I$, and such that $F$ and $G$ are continuous, and differentiable. Then there exists an equilibrium $\tau^{*}\in I$ which is locally stable, while the equilibria $\pm\infty$ are unstable.
@@
We know already from Theorem 6 that at least one equilibrium exists in the interval $I$. Again, we let H(x)=G(x)F(x)H(x)=G(x)-F(x). Note first that since $H(a)>0$ and $H(b)<0$, by continuity and differentiability $H$ must cross 0 at some point $\tau_{0}$ with $H^{\prime}(\tau_{0})<0$. Assume not: then $H(x)\geq 0$ for all $x$, which contradicts the fact that $H(x)<0$ for $x\geq b$. It follows by Theorem 5 that $\tau_{0}$ is a stable equilibrium.
@@
    Pick the equilibrium $\tau_{2}$ which is closest to $b$ (this could be the same as the $\tau_{1}$ we chose in the previous step). By Condition 2 and since $H(x)$ is continuous, $H(x)<0$ for $x>\tau_{2}$. It follows that $H^{\prime}(\tau_{2})\leq 0$. Then $\tau_{2}$ is stable, at least on the right (if $H^{\prime}(\tau_{2})=0$, then $\tau_{2}$ is stable on the right and unstable on the left).
We know by Condition 2 that $H(x)\neq 0$ for $x\notin[a,b]$, so that no equilibria occur outside the interval $I$ other than $\pm\infty$. Since both $\tau_{1}$ and $\tau_{2}$ (or just $\tau_{1}$, if they are the same equilibrium) are stable (at least to the left for $\tau_{1}$, and right for $\tau_{2}$), it is topologically necessary that the uninformative equilibria $\pm\infty$ are unstable. ∎
@@
Appendix C Omitted Results for RBTS
In this section we formally analyze the RBTS mechanism under real-valued signals; fully characterize equilibria under monotonicity and continuity conditions; and derive the dynamics of equilibria under the Gaussian model. We focus on RBTS instead of its predecessor, the Bayesian Truth Serum (BTS), because incentive alignment under BTS requires an arbitrarily large number of agents; meanwhile, RBTS guarantees truthfulness for $n\geq 3$ agents.
C.1 Equilibrium characterization
As a reminder, in RBTS each agent $i$ submits both an information report $r_{i}\in\{L,H\}$ and also a prediction report $p_{i}\in[0,1]$. Let $S(p,r)=2p\mathbf{1}[r=H]-p^{2}$. RBTS randomly picks a reference agent $j$ and peer agent $k$, and pays agent $i$
$M_{\textrm{RBTS}}((r_{i},p_{i}),(r_{j},p_{j}),(r_{k},p_{k}))=S(p_{j}+\delta t(r_{i}),r_{k})+S(p_{i},r_{k}),$
where t(L)=1t(L)=-1, $t(H)=1$, and $\delta>0$ is chosen by the mechanism. We refer to the first term in Equation 10 as the information score, and the second as the prediction score. Under binary signals and reports, truthfulness over both the information and prediction report forms a strict Bayes-Nash equilibrium.
Proposition 9.
$\tau^{*}=\pm\infty$ are threshold equilibria under RBTS.
Proof.
We also observe necessary and sufficient conditions for finite equilibria $\tau$, relative to the function $G$. Let $Q(x)=\mathop{\mathbb{E}}_{x^{\prime}\sim\beta(x)}P(x;x^{\prime})$, and let $\bar{P}(\tau;x)=\mathop{\mathbb{E}}_{x^{\prime}\sim\beta(x)}P(\tau;x^{\prime})$ be the expected conditional probability of another agent reporting $L$ over signal $x$.